# Addressing Hallucinations with RAG and NMISS in Italian Healthcare LLM Chatbots

### v2

@@ Abstract
I combine detection and mitigation techniques to addresses hallucinations in Large Language Models (LLMs). Mitigation is achieved in a question-answering Retrieval-Augmented Generation (RAG) framework while detection is obtained by introducing the Negative Missing Information Scoring System (NMISS), which accounts for contextual relevance in responses. While RAG mitigates hallucinations by grounding answers in external data, NMISS refines the evaluation by identifying cases where traditional metrics incorrectly flag contextually accurate responses as hallucinations. I use Italian health news articles as context to evaluate LLM performance. Results show that Gemma2 and GPT-4 outperform the other models, with GPT-4 producing answers closely aligned with reference responses. Mid-tier models, such as Llama2, Llama3, and Mistral benefit significantly from NMISS, highlighting their ability to provide richer contextual information. This combined approach offers new insights into the reduction and more accurate assessment of hallucinations in LLMs, with applications in real-world healthcare tasks and other domains.
@@ 1 Introduction
Large Language Models (LLMs) are AI systems typically built on the Transformer architecture, as introduced by Vaswani [52]. Designed to understand and generate human language, LLMs are trained on extensive text datasets, allowing them to grasp the nuances of language. These models perform various tasks, from basic text classification to advanced text generation, and more recent iterations can even interpret images and generate code.
@@
LLMs have surged in popularity due to their intuitive, prompt-based interfaces, enabling effortless interaction. The past few years have seen the continuous development of notable LLMs like ChatGPT [33, 1], Gemini [11], Falcon [37], Llama [49, 50, 30], and Gemma [46, 47]. As highlighted by [15], these models excel in language comprehension [14, 16], text summarization [59], and reasoning [54, 22, 38]. As a result, their applications span various industries, such as processing electronic medical records, matching clinical trials, and drug discovery in healthcare. In finance, LLMs assist with fraud detection and trading, while in business, they automate customer service via chatbots and virtual assistants [35].
@@
However, LLMs face a significant challenge: hallucinations [12, 51]. Hallucinations occur when models generate nonsensical or inaccurate content, failing to reflect the source material accurately [17]. This behavior is a serious threat since it affects the application of LLMs to real-world cases [20], as hallucinations can undermine the reliability of their outputs. [15] classifies hallucinations into two types: factuality hallucinations, where generated content diverges from verifiable facts, and faithfulness hallucinations, where the content strays from user instructions or input context. Factuality hallucinations manifest as either factual inconsistencies or fabrications. For example, a model might mistakenly claim that “The Colosseum is located in Berlin, Germany” instead of Rome or invent details about non-existent technologies. Faithfulness hallucinations occur when the output deviates from user instructions or context. These can involve instruction inconsistencies (e.g., responding in Spanish when asked to translate into Italian), context inconsistencies (introducing unrelated information), or logical errors (e.g., miscalculating “two times four” as three).
@@
The mitigation of hallucinations has become an urgent concern due to the widespread use of LLMs in sensitive areas such as healthcare. Addressing hallucinations is challenging due to LLMs’ reliance on reinforcement learning from human feedback [34], their opaque black-box nature that hides knowledge boundaries [45, 42], and their training on large datasets that may contain outdated or incorrect information [60]. Additionally, their versatility across many tasks complicates mitigation efforts, while their strong performance can make hallucinations, especially factual fabrications, harder to detect.
@@
Researchers focused on various techniques to mitigate hallucinations. One of the most promising methods is Retrieval-Augmented Generation (RAG) [24], which incorporates external knowledge during text generation to anchor outputs in factual information. RAG addresses issues of accuracy and currency in LLM responses by using authoritative sources, generating verifiable answers, and enhancing user confidence across applications [21]. Other approaches include Prompt Engineering that involves optimizing LLMs outputs with different instructions [48, 55], self-refinement through feedback and reasoning [27], where feedback on LLM outputs improves the accuracy and quality of subsequent responses, and prompt tuning, which modifies instructions during fine-tuning to improve task-specific performance [23].
@@
Mitigating hallucinations requires first their detection, and various metrics have emerged. Traditional metrics like ROUGE [26], BLEU [36], and METEOR [4] measure consistency between generated text and source text, but they often struggle with semantically relevant but superficially divergent text [5]. To address these limitations, newer embedding-based methods such as BERTScore [58] use contextual embeddings to assess semantic similarity between sentences. Another approach involves ”opening the box” where token probabilities or logit outputs of LLMs are analyzed to detect hallucinations [3, 51]. However, these techniques require access to internal model states, which is not always feasible for proprietary systems.
This paper proposes a general framework for detecting and mitigating hallucinations, with an application to the Italian healthcare domain. For the mitigation part, I employ a RAG architecture to integrate external knowledge into the generation process, ensuring that the models are grounded in factual data. The methodology applies a multilingual Sentence Transformer model [40, 41] for vector database generation and retrieval. I compare a range of LLMs, including Gemma [46], Gemma 2 [47], Mistral [18], GPT-4 [1], Llama2 [50], and Llama3 [30] for the answer generation. This framework uses a prompting strategy to instruct models to generate answers based on the provided context. To assess interpretability and difficulty, I categorize 100 questions into three levels of understanding: low, mid, and high. For the detection strategy, I introduce the Negative Missing Information Scoring System (NMISS), a novel scoring framework designed to be applied alongside existing metrics to integrate contextual relevance. By analyzing residual words from machine-generated answers after comparison with human responses, NMISS refines traditional metrics by highlighting cases where they may unfairly penalize performance despite the enhanced contextual relevance of the responses.
Results show that Gemma2 and GPT-4 achieve the highest overall performance in Italian healthcare-related question-answering tasks across all difficulty levels, while Gemma exhibits the highest rate of hallucinations. The NMISS scoring system provides a more accurate assessment of LLM performance, especially in high-difficulty tasks for mid-tier models like Llama2, Llama3, and Mistral, which show notable improvements compared to traditional metrics. Additionally, NMISS recognizes the value of contextual information provided by models in cases where they fail to deliver exact matches to expected answers in difficult tasks. By focusing on the relevance and coherence of the provided context, NMISS highlights its advantage in evaluating responses that are partially correct or contextually appropriate, even when they do not fully align with the reference answers.
@@
This paper is organized as follows: Section 2 reviews the relevant literature on hallucination detection and mitigation. Section 3 details the methodology, highlighting the contributions of NMISS. Section 4 describes the experimental setup on Italian healthcare news data, and Section 5 presents my findings. Finally, Section 6 summarizes contributions and outlines future research directions.
@@ 2.1 Hallucination Detection Techniques
Hallucinations in LLMs have been widely studied. [15] provides a detailed overview of strategies to detect and mitigate hallucinations, offering insight into their occurrence in current LLMs architectures. Ensuring LLM’s trustworthiness depends on early detection, as emphasized by [51], who highlights the importance of proactive measures during text generation.
[60] categorize detection methods into three main types. First, logit-based estimation, which uses token-level probabilities or entropy from the model’s logits [39]. Second, verbalized confidence estimation, where LLMs are asked to self-rate the confidence in their responses. Third, consistency-based estimation involves generating multiple candidate outputs and comparing them to detect logical inconsistencies, mimicking human reasoning to refine answers [53, 44, 61].
Furthermore, [19] categorize hallucination detection into white-box, grey-box, and black-box techniques. White-box methods analyze the model’s internal workings to identify when incorrect statements are generated. Grey-box approaches examine logit outputs to find areas of low confidence in the output. Black-box techniques, which only require the model’s final answers, are simpler to apply since they don’t require access to internal states.
Recent studies introduce SAPLMA [3], which detects hallucinations by using a classifier trained to predict the truthfulness of statements based on LLMs’ hidden layer activations. This approach outperforms traditional methods that rely solely on sentence probabilities. [51] propose an iterative process that detects potential hallucinations by identifying tokens with low normalized probabilities. The system then retrieves relevant information to correct these errors, effectively reducing hallucinations in GPT-3.5.
@@ 2.2 Resource-Intensive Mitigation Strategies
In addition to detection, various strategies have been developed to mitigate hallucinations in LLMs. One common approach is Knowledge Augmentation, such as Knowledge Injection (KI). [29] enhance prompts with information from knowledge graphs, such as adding details like location names for review tasks. This significantly reduces hallucinations and improves factual accuracy, with KI-LLMs producing more correct assertions and outperforming larger models in response quality.
[31] take a direct approach by training T5 models on knowledge graph triples from Wikidata. By masking key entities, the model learns and generalizes knowledge effectively. SKILL, after fine-tuning, outperforms standard T5 models on question-answering tasks, proving the effectiveness of structured data infusion.
Another key mitigation strategy is retrieval-based frameworks. [25] combines a sequence-to-sequence model with a dense vector index of Wikipedia articles using the RAG framework. This integration enhances the accuracy and reliability of generated content. Similarly, [13] apply retrieval techniques in biomedical contexts, incorporating sources like UMLS and Wikipedia to generate more factual summaries, which are crucial for healthcare applications.
[62] introduce a self-supervision framework that assigns risk scores, called POLAR scores, to LLM responses as indicators of hallucination risk. Their approach uses Pareto optimization to balance supervision from multiple sources. The process involves gathering LLM responses, training a probabilistic function $h$ to estimate error rates, and optionally re-prompting the LLM based on the POLAR score with RAG support. Experimental results show significantly reduced hallucination rates across tasks like question answering and text generation.
[57] propose the MixAlign framework, which aligns user queries with external knowledge to resolve ambiguities. This improves response accuracy and reduces hallucinations, with results showing a sharp increase in correct answers and a decrease in hallucinations.
@@ 2.3 Zero-Resource Hallucination Strategies
The strategies discussed in Section 2.2 are resource-intensive. In contrast, zero-resource hallucination methods are designed for black-box models and scenarios with limited resources, relying only on model outputs without needing internal access or external data. Many of these techniques not only detect hallucinations but also help mitigate them by refining responses based on internal consistency assessments.
@@
One approach is through self-consistency techniques, which compare multiple responses generated by the LLM itself. [28] propose a zero-resource method that generates multiple responses from GPT-3 and compares them to a reference answer to assess factual consistency. This method detects and mitigate hallucinations using low-level metrics and outperforms traditional grey-box and black-box methods.
@@
Another approach is self-rating techniques, as proposed by [6]. The authors introduce RelD, a discriminator designed to detect hallucinations in LLM-generated answers. RelD is trained by converting a regression task into a binary classification task, aligning with human intuition. By flagging low-confidence responses, it serves as an indirect mitigation mechanism. Experiments across various datasets and models (e.g., BLOOM [56]) show RelD’s effectiveness, validated by both automatic metrics and human-in-the-loop evaluations. Similarly, [2] addresses hallucinations in fabricated references using Direct Queries (DQs) and Indirect Queries (IQs) without external resources. DQs involve asking the model a binary question multiple times to calculate a probability. In contrast, IQs ask about the authors or content of the reference and check for consistency across responses. Although primarily designed for detection, these methods can also reduce hallucinations by discarding low-confidence responses. Results show that IQs generally perform better than DQs in detecting hallucinations, though performance varies among models like GPT-4, GPT-3, and Llama 2.
@@
Finally, [32] explore self-contradiction techniques as a factor in hallucinations. They introduce a prompting framework to detect and mitigate contradictions within the same response, relying solely on the model’s logical reasoning capabilities. This framework includes functions for gathering contextual information, generating sentences, identifying contradictions, and refining sentences. Across various LLMs, this approach achieves an 80% F1 score.
2.4 Metrics for Evaluating Hallucinations
Traditional metrics to detect or manage hallucinations [5] are notably ROUGE [26], BLEU [36], and METEOR [4]. These indicators measure the consistency between the generated text and the source text. However, a drawback of these simple metrics is that they do not accurately capture distance word-dependencies or if the two answers diverge in lexicon but have equal meaning. BERTScore overcomes this issue [58] by leveraging pre-trained contextual embeddings from BERT [9] and matches words in candidate and reference sentences by cosine similarity, correlating well with human judgments.
2.5 Contributions
Despite significant progress in hallucination detection and mitigation, key challenges persist. Many existing methods, such as knowledge injection or retrieval, are resource-intensive, limiting their practicality in certain applications. Additionally, traditional evaluation metrics often fail to account for contextual accuracy, unfairly penalizing semantically correct responses that differ lexically from reference answers.
This paper addresses these gaps in several steps. The key contributions of this work are as follows:
(1) A combined approach for hallucination detection and mitigation in LLMs. I develop a RAG-based architecture to integrate external knowledge into the generation process, ensuring that models are grounded in factual data. The methodology leverages a multilingual Sentence Transformer model [40, 41] for vector database generation and retrieval. This framework includes a prompting strategy to instruct models to generate answers based on the provided context, enabling a comparative evaluation of various LLMs using this approach.
(2) Comprehensive analysis of hallucinations across reasoning levels. To evaluate hallucination tendencies in LLMs, I analyze their performance across progressively challenging levels of reasoning, categorized in low, mid and high levels. This approach enables an assessment of how hallucination rates and mitigation effectiveness vary as tasks demand increasing degrees of logical complexity and understanding. The progressive evaluation highlights how model reliability and factual grounding are affected under escalating reasoning challenges, providing insights into their strengths and vulnerabilities.
(3) Development of a detection scoring system. I introduce the Negative Missing Information Scoring System (NMISS), a novel framework that fine-grains traditional metrics by incorporating contextual relevance. By analyzing residual words from machine-generated answers after comparison with human responses, NMISS highlights cases where traditional metrics may penalize models despite the enhanced contextual relevance of their answers.

### v3

@@ Abstract
This work addresses hallucinations in Large Language Models (LLMs) by combining detection and mitigation strategies within a knowledge-driven evaluation framework. I employ a Retrieval-Augmented Generation (RAG) setup to ground model responses in structured external data, thereby improving factual reliability and helping models better follow instructions in zero-shot prompting settings. To enhance evaluation beyond surface-level overlap, I introduce the Negative Missing Information Scoring System (NMISS), a novel scoring system that complements traditional metrics such as BLEU, ROUGE, and METEOR by incorporating contextual relevance. NMISS helps distinguish between genuine hallucinations and responses that correctly infer missing information from trusted sources. Applied to a health-related question-answering task using Italian news articles, NMISS reveals important differences in model behavior. While Gemma2 and GPT-4 produce the most reference-aligned answers under standard metrics, Gemma2 also achieves the highest NMISS outperformance, suggesting that its semantically rich outputs are often under-recognized by traditional scores. In contrast, GPT-4 gains less from NMISS due to its strict adherence to prompt phrasing, leaving little room for contextual enhancement. Mid-tier models such as LLaMA2, LLaMA3, and Mistral show substantial NMISS gains when scored with BLEU and ROUGE-1, indicating their ability to provide contextually meaningful content even when traditional scores are lower. Overall, the RAG–NMISS framework improves hallucination detection, reveals context-aware model behavior often overlooked by traditional metrics, and supports more reliable evaluation in real-world structured knowledge tasks.
@@ 1 Introduction
Large Language Models (LLMs) are AI systems typically built on the Transformer architecture, as introduced by vaswani2017attention. Designed to understand and generate human language, LLMs are trained on extensive text datasets, allowing them to grasp the features of language. These models perform various tasks, from basic text classification to advanced text generation, and more recent iterations can even interpret images and generate code.
@@
LLMs have surged in popularity due to their intuitive, prompt-based interfaces, enabling effortless interaction. The past few years have seen the continuous development of notable LLMs like ChatGPT [OpenAI2022, achiam2023gpt], Gemini [Google2023], Falcon [penedo2023refinedweb], Llama [touvron2023llama, touvron2023llama2, meta2024llama3], and Gemma [team2024gemma, team2024gemma2]. As highlighted by huang2023survey, these models excel in language comprehension [hendrycks2020measuring, huang2024c], text summarization [zhang2024benchmarking], and reasoning [wei2023simple, kojima2022large, qiao2022reasoning]. As a result, their applications span various industries, such as processing electronic medical records, matching clinical trials, and drug discovery in healthcare. In finance, LLMs assist with fraud detection and trading, while in business, they automate customer service via chatbots and virtual assistants [ozdemir2023quick].
@@
However, LLMs face a significant challenge: hallucinations [guerreiro2023hallucinations, varshneystitch]. Hallucinations occur when models generate nonsensical or inaccurate content, failing to reflect the source material accurately [ji2023survey]. This behavior is a serious threat since it affects the application of LLMs to real-world cases [kaddour2023challenges], as hallucinations can undermine the reliability of their outputs. [huang2023survey] classifies hallucinations into two types: factuality hallucinations, where generated content diverges from verifiable facts, and faithfulness hallucinations, where the content strays from user instructions or input context. Factuality hallucinations manifest as either factual inconsistencies or fabrications. For example, a model might mistakenly claim that “The Colosseum is located in Berlin, Germany” instead of Rome or invent details about non-existent technologies. Faithfulness hallucinations occur when the output deviates from user instructions or context. These can involve instruction inconsistencies (e.g., responding in Spanish when asked to translate into Italian), context inconsistencies (introducing unrelated information), or logical errors (e.g., miscalculating “two times four” as three).
@@
The mitigation of hallucinations has become an urgent concern due to the widespread use of LLMs in sensitive areas such as healthcare. Addressing hallucinations is challenging due to LLMs’ reliance on reinforcement learning from human feedback [ouyang2022training], their opaque black-box nature that hides knowledge boundaries [sun2023moss, ren2023investigating], and their training on large datasets that may contain outdated or incorrect information [zhang2023siren]. Additionally, their versatility across many tasks complicates mitigation efforts, while their performance makes hallucinations, especially factual fabrications, harder to detect.
@@
Researchers have explored several strategies to mitigate hallucinations. One promising approach is Retrieval-Augmented Generation (RAG) [lewis2020rag], which incorporates external knowledge during generation to ground outputs in factual information. RAG improves both the accuracy and timeliness of model responses by leveraging authoritative sources, producing verifiable answers, and increasing user confidence [kang2023ever]. Other techniques include prompt engineering, which optimizes outputs via instruction design [tonmoy2024comprehensive, white2023prompt]; self-refinement, where feedback on prior outputs enhances subsequent responses [madaan2023selfrefineiterativerefinementselffeedback]; and prompt tuning, which adapts instruction embeddings during fine-tuning to improve task-specific performance [lester-etal-2021-power].
@@
Detecting hallucinations is a prerequisite for effective mitigation. Traditional metrics such as ROUGE [lin2004rouge], BLEU [papineni2002bleu], and METEOR [banerjee2005meteor] measure lexical overlap between generated and reference text but often fail to capture semantic equivalence [celikyilmaz2020evaluation]. More recent embedding-based metrics, such as BERTScore [zhang2019bertscore], use contextual embeddings to assess semantic similarity. Another direction involves “opening the box”—analyzing token probabilities or logit outputs to detect hallucinations [azaria2023internal, varshneystitch]—though this approach requires access to internal model states, which is not feasible for proprietary systems.
This work proposes a general, domain-independent framework for hallucination detection and mitigation in LLMs. For mitigation, a RAG architecture integrates external knowledge into the generation process, ensuring that outputs remain grounded in factual information. The methodology employs a multilingual Sentence Transformer model [reimers-2019-sentence-bert, reimers-2020-multilingual-sentence-bert] to construct and query a vector database for contextual retrieval. A range of LLMs, specifically Gemma [team2024gemma], Gemma-2 [team2024gemma2], Mistral [jiang2023mistral], GPT-4 [achiam2023gpt], Llama-2 [touvron2023llama2], and Llama-3 [meta2024llama3] are compared using a zero-shot prompting strategy that instructs models to generate answers based on retrieved context. To evaluate interpretability and task complexity, 100 questions are categorized into three difficulty levels: low, medium, and high.
For detection, this work introduces the Negative Missing Information Scoring System (NMISS), a novel evaluation layer that extends traditional metrics by incorporating contextual relevance. Unlike prior metrics, which assess only lexical or semantic overlap with reference answers, NMISS leverages context from external sources to distinguish genuine hallucinations from contextually valid elaborations. Specifically, it examines lexical items present in model outputs but absent from human references, evaluating whether these additions are substantiated by the provided context. Through this mechanism, NMISS refines traditional scoring methods to more accurately reflect whether supplementary content contributes relevant information or constitutes hallucination.
Under the RAG-based evaluation, Gemma-2 and GPT-4 achieve the highest overall performance across all difficulty levels, demonstrating strong grounding and retrieval effectiveness, while Gemma exhibits the highest hallucination rate. NMISS, in turn, provides a more reliable assessment of model behavior, particularly in high-difficulty cases for mid-tier models such as Llama-2, Llama-3, and Mistral, which show notable improvements under this context-aware metric. Interestingly, models with more verbose generation styles tend to benefit from NMISS, as their outputs often retrieve and reuse lexical elements directly from the provided context. This might reflect computational constraints or limited abstraction capacity. Nevertheless, Gemma-2 also benefits from this behavior, suggesting that contextual retrieval contributes positively to factual grounding rather than mere verbosity. By emphasizing contextual relevance and coherence over strict lexical matching, NMISS highlights its advantage in evaluating responses that are partially correct or contextually appropriate, even when they diverge from the reference answers.
The main contributions of this paper are as follows: (i) it proposes NMISS, a novel scoring framework for hallucination detection that extends standard evaluation metrics by incorporating contextual relevance, thereby enabling a more accurate distinction between hallucinations and contextually valid elaborations. The framework is general-purpose and applicable beyond the healthcare domain; (ii) it develops a RAG–based evaluation framework that grounds LLM outputs in structured external knowledge, demonstrating that retrieval mitigates hallucinations and enhances factual grounding; (iii) it presents a comprehensive evaluation of multiple LLMs across varying levels of reasoning difficulty, showing that NMISS exposes performance differences overlooked by traditional metrics; and (iv) it introduces a new dataset comprising 100 Italian healthcare questions annotated by reasoning complexity, supporting domain-specific benchmarking and multilingual question-answering evaluation.
@@
The work is organized into five sections. Section 2 reviews the relevant literature on hallucination detection and mitigation. Section 3 presents the proposed methodology and highlights the contributions of NMISS. Section 4 describes the experimental setup on Italian healthcare news data, and Section 5 reports the findings. Finally, Section 6 concludes the study by summarizing the main contributions and outlining future research directions.
@@ 2 Related Works
This section reviews recent research on hallucination management in LLMs, organized along two main dimensions. First, it distinguishes between detection techniques, which aim to identify hallucinated content, and mitigation techniques, which seek to reduce hallucinations during text generation. Second, within mitigation, existing approaches are further classified by their resource requirements: resource-intensive strategies that rely on model internals or external data, and zero-resource strategies that operate in black-box settings and often combine elements of both detection and mitigation. This organization clarifies the conceptual landscape of hallucination research and situates the contributions of this work within it.
@@ 2.1 Hallucination Detection Techniques
Hallucinations in LLMs have been widely studied in recent years. huang2023survey provide a comprehensive overview of strategies for detecting and mitigating hallucinations, offering valuable insight into how they emerge in current LLM architectures. Ensuring the trustworthiness of LLMs depends critically on early detection, as emphasized by varshneystitch, who highlight the importance of proactive identification during text generation.
zhang2023siren categorize detection methods into three main types. First, logit-based estimation, which relies on token-level probabilities or entropy derived from model logits [quevedo2024detecting]. Second, verbalized confidence estimation, in which LLMs are prompted to self-rate their confidence in generated responses. Third, consistency-based estimation, which compares multiple generated outputs to identify logical inconsistencies, thereby mimicking human reasoning to refine responses [wang2022self, shi2023replug, zhao2023verify].
Complementary to this taxonomy, jiang2024hallucination classify hallucination detection methods into white-box, grey-box, and black-box approaches. White-box methods analyze internal model activations to pinpoint when incorrect statements are produced, while grey-box methods examine logit-level confidence to detect uncertainty. In contrast, black-box methods rely solely on model outputs and are easier to apply, since they do not require access to internal states.
Recent studies have also introduced specialized architectures. SAPLMA [azaria2023internal], for example, detects hallucinations using a classifier trained on hidden layer activations to predict the truthfulness of statements, outperforming traditional probability-based baselines. Likewise, varshneystitch propose an iterative detection process that identifies low-probability tokens and retrieves external evidence to correct potential errors, effectively reducing hallucinations.
Alongside model-internal techniques, several surface-level metrics have been employed to detect hallucinations based on output similarity. Traditional reference-based metrics such as ROUGE [lin2004rouge], BLEU [papineni2002bleu], and METEOR [banerjee2005meteor] measure lexical overlap between generated and reference text, but often fail to capture semantically valid variations or long-range dependencies [celikyilmaz2020evaluation]. BERTScore [zhang2019bertscore], which leverages contextual embeddings from BERT [devlin2018bert], improves upon these by evaluating token-level semantic similarity. However, even BERTScore does not account for whether additional, unmatched content is contextually justified.
This study differs from prior work by introducing the Negative Missing Information Scoring System (NMISS), a context-aware scoring framework that enhances existing metrics through contextual alignment, without requiring access to model internals or probability distributions.
@@ 2.2 Resource-Intensive Mitigation Strategies
Beyond detection, a range of strategies have been developed to mitigate hallucinations in LLMs, often requiring computational or data resources. One prominent category is Knowledge Augmentation, including techniques such as Knowledge Injection (KI). martino2023knowledge enhance prompts with information derived from knowledge graphs, such as adding geographical details for review tasks. This approach significantly reduces hallucinations and improves factual accuracy, with KI-enhanced LLMs producing more correct assertions and outperforming larger baseline models in response quality. moiseev2022skill adopt a more direct approach by training T5 models on Wikidata triple knowledge graphs. By masking key entities during training, the model learns to infer and generalize structured knowledge effectively. Their model outperforms standard T5 variants on question-answering benchmarks, demonstrating the benefits of structured data infusion.
Another major line of research involves retrieval-based frameworks. lewis2021improving combine a sequence-to-sequence model with a dense vector index of Wikipedia articles using the RAG framework, thereby improving the accuracy and factual grounding of generated content. Similarly, guo2024retrieval apply retrieval augmentation in biomedical contexts, using external sources to produce more accurate and verifiable summaries, which are crucial for healthcare applications.
A different approach, introduced by zhao2023automatic, is a self-supervision framework that assigns risk scores, termed POLAR scores, to LLM responses as indicators of hallucination likelihood. Their method employs Pareto optimization to balance supervision signals from multiple sources. The process involves collecting model responses, training a probabilistic function to estimate error rates, and optionally re-prompting the LLM based on the POLAR score with RAG support. Experimental results demonstrate substantial reductions in hallucination rates across various tasks, including question answering and text generation.
In a related effort, zhang2024knowledgealignmentproblembridging propose the MixAlign framework, which aligns user queries with external knowledge sources to resolve semantic ambiguities. This method improves both response accuracy and factual consistency, yielding a marked increase in correct answers and a reduction in hallucination frequency.
Unlike prior resource-intensive frameworks, this study introduces a lightweight mitigation approach that integrates zero-shot prompt engineering with RAG for factual grounding in the healthcare domain.
@@ 2.3 Zero-Resource Hallucination Strategies
The strategies discussed in Section 2.2 are resource-intensive. In contrast, zero-resource hallucination methods are designed for black-box models and scenarios with limited computational capacity, relying solely on model outputs without requiring internal access or external training data. Many of these techniques not only detect hallucinations but also mitigate them by refining responses through internal consistency checks.
@@
One common approach involves self-consistency techniques, which compare multiple responses generated by the LLM itself. manakul2023selfcheckgpt propose a zero-resource method that samples multiple outputs from GPT-3 and compares them to a reference answer to assess factual consistency. This method detects and mitigates hallucinations using low-level statistical metrics and outperforms traditional grey-box and black-box baselines.
@@
Another class of methods includes self-rating techniques, as proposed by chen2023hallucination. The authors introduce RelD, a discriminator designed to detect hallucinations in LLM-generated answers by converting a regression task into a binary classification problem, aligning with human judgment. By flagging low-confidence responses, it serves as an indirect mitigation mechanism. Experiments across several datasets and models (e.g., BLOOM [workshop2022bloom]) demonstrate RelD’s effectiveness, validated through both automated metrics and human evaluations. Similarly, agrawal2023language address hallucinations in fabricated references using Direct Queries (DQs) and Indirect Queries (IQs) without external resources. DQs repeatedly ask binary questions to estimate response probability, whereas IQs query contextual details such as authors or publication content to test internal consistency. Although primarily designed for detection, these approaches can also mitigate hallucinations by filtering out low-confidence responses. Results show that IQs generally outperform DQs, though effectiveness varies among models such as GPT-4, GPT-3, and Llama-2.
@@
Finally, mundler2023self explore self-contradiction techniques, which identify hallucinations by detecting logical inconsistencies within a single response. Their prompting framework relies entirely on the model’s internal reasoning capabilities and includes procedures for gathering contextual information, generating candidate sentences, identifying contradictions, and refining responses. Across several LLMs, this method achieves an F1 score of approximately 80%.
In contrast to self-referential, zero-resource techniques, this work introduces a context-aware hallucination detection mechanism, NMISS, which evaluates generated outputs without relying on self-consistency signals or model confidence estimation. The proposed hybrid approach remains lightweight, requiring neither fine-tuning nor model access, while enhancing factual grounding through the integration of retrieved evidence.